OpenStudy (imtiaz7):

Prove that the rate of heat production in each of the two resistors connected in parallel are in versely proportional to the resistances.

OpenStudy (binary3i):

rate of heat = I^2R where I=V/R so it becomes V^2/R so you see inversely proportional